This is my problem:
Find the coordinates of any points of intersection.

3x^2 + 7y^2 = 187
3x^2 - 7y = 47
I attempted to change the equations to all y =
But not having any luck with what I end up and then attempting to graph them to find the intersections.
Thank you!

Answer by Alan3354(69443)

Find the coordinates of any points of intersection.
3x^2 + 7y^2 = 187 eqn 1
3x^2 - 7y = 47 eqn 2
eqn 1 is a circle, and eqn 2 is a parabola. There can be a max of 4 intersections.
I don't know how to do the graphs on this site, tho I have seen others do them.
-------------------------
-------------------
Subtract (I didn't see it at first, I was working on substition)
0x^2 +7y^2 +7y = 140
y^2 + y - 20 = 0
(y+5)*(y-4) = 0
y = 4, y = -5
-----------
Sub y = 4 into eqn 2
3x^2 -28 = 47
x^2 = 25
x = +5, -5
----------
Sub y = -5 into eqn 2
3x^2 +35 = 47
x^2 = 4
x = +2, -2
----------------
So the intersections are:
(2,-5), (-2,5) and (5,4), (5,-4)
I started with substitution, which would have been a lot of work, but the answers would be the same.